OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x f(x) = x-8/x+7 and g(x) = -7x-8/x-1

OpenStudy (anonymous):

f(g(x)) = f (-7x-8/x-1) Subs this value of x in the equation (f(x) = x-8/x+7)and see ...

OpenStudy (anonymous):

What??

OpenStudy (anonymous):

An inverse is when you replace X with Y... For example, the inverse of y = 3x is x = 3y (Thus x/3 = y) For this question, find the inverse of f(x), and your answer should equal g(x).

OpenStudy (anonymous):

f(x) = x-8/x+7 f((-7x-8/x-1) = [ (-7x-8)/(x-1) - 8 ] / [ (-7x-8)/(x-1) + 7 ] = -15x/-15 = x

OpenStudy (anonymous):

And Proceed Same as this for g(f(x))

OpenStudy (anonymous):

your job is to compute \[f\circ g(x)\] first step is \[f\circ g(x)=f(g(x))=f(\frac{-7x-8}{x-1})\] \[=\frac{\frac{-7x-8}{x-1}-8}{\frac{-7x-8}{x-1}+7}\] then simplify this ugly compound fraction start by multiplying to and bottom by \(x-1\) you will have a whole orgy of cancellation, leaving nothing but \(x\)

OpenStudy (anonymous):

can you please explain to me how to do the g(f(x) because i am not seeing how i can plug it in to prove it?

OpenStudy (anonymous):

\[g(f(x))=g(\frac{x-8}{x+7})\] is as start

OpenStudy (anonymous):

\[g(\frac{x-8}{x+7})=\frac{-7\left(\frac{x-8}{x+7}\right)-8}{\left(\frac{x-8}{x+7}\right)-1}\]

OpenStudy (anonymous):

that is, replace the input by \(\frac{x-8}{x+7}\) now it is algebra from here on in ugly algebra, but algebra just the same

OpenStudy (anonymous):

thank you so much!!